News by Date
News by Category
Solar System
White Dwarfs
Neutron Stars
Black Holes
Milky Way Galaxy
Normal Galaxies
Galaxy Clusters
Cosmology/Deep Field
Press Resources
Status Reports
Press Advisories
Image Releases
Release Guidelines
Image Use Policy
Web Shortcuts
Chandra Blog
RSS Feed
Email Newsletter
News & Noteworthy
Image Use Policy
Questions & Answers
Glossary of Terms
Download Guide
Get Adobe Reader
Chandra @ NASA
Visit the Chandra pages at the NASA portal (opens in new window)
Image Use
Image Use Policy & Request Form
Guidelines for utilizing images, applets, movies, and animations featured in this Web Site.
Getting Hard Copies of Images
Ways to obtain photos, slides, etc of Chandra images.
Operations CXO Status Report

Friday 8 April 2005 9.00am EST

During the last week the observing schedule was interrupted following a high radiation event at 8:59pm on Apr 5 that stopped the daily load and activated the science instrument safing sequence (SCS 107). The event occurred just prior to radiation belt entry and is thought to be due to a 'puffed-up' belt state caused by high levels of coronal stream activity. The loads were replanned and science was re-started at radiation belt exit on Apr 6 with 2.7 Ks of science loss. As a result of the SCS 107 run the attitude during perigee was not as planned and resulted in an elevated temperature for the IEPHIN causing a rail voltage drop. This was anticipated in the re-plan and considered acceptable for this case.

A Chandra press release was issued on Apr 6 describing deep X-ray observations that combined with sub-mm data provide evidence that some extremely luminous adolescent galaxies and their central black holes underwent a phenomenal spurt of growth more than 10 billion years ago. A second release was issued on Apr 7 describes a study of "fossil galaxies" - ancient galaxy groups in which all of the large galaxies have gradually merged to form one central giant galaxy. For details see:

The schedule of targets for the next week is shown below and includes a DDT observation of HESS J1826-148 coordinated with HESS.

NGC5353                ACIS-I       Apr 10
SDSS J115314.36+03     ACIS-S
IRAS05436-0007         ACIS-I
Radiation Belts                     Apr 11
E0102-72               ACIS-S
Cas A                  ACIS-I       Apr 12
Cas A                  ACIS-S
Cas A                  HRC-I
FIRST J131739.2+41     ACIS-S
RCS1107.3-0523         ACIS-S
HESS J1826-148         ACIS-S
A2142                  ACIS-I       Apr 13
2QZ J114327.3-0050     ACIS-S
Radiation Belts
87GB 1143+4437         ACIS-S       Apr 14
3C 325                 ACIS-S
NGC 5548               HRC-S/LETG   Apr 15
Radiation Belts                     Apr 16
3C 325                 ACIS-S       Apr 17
NGC 5548               HRC-S/LETG
E0102-72               ACIS-S


All spacecraft subsystems continued to support nominal operations.

Return to Status Reports